### v4

# Non-Resolution Reasoning (NRR): A Computational Framework for Contextual Identity and Ambiguity Preservation

Abstract
Current artificial intelligence systems, despite remarkable capabilities in text generation and pattern recognition, exhibit a fundamental architectural limitation: they resolve ambiguity prematurely. This premature semantic collapse—the tendency to collapse multiple valid interpretations into a single output—stems from classical identity assumptions embedded in standard neural architectures. We propose Non-Resolution Reasoning (NRR), a computational framework that treats ambiguity retention as a valid reasoning mode rather than a defect to be eliminated. NRR introduces three core principles: (1) Non-Identity ($A\neq A$)—the same symbol refers to different entities across contexts; (2) Approximate Identity ($A\approx A$)—entities share partial structural overlap without being identical; and (3) Non-Resolution—conflicting interpretations can coexist without forced convergence. We formalize these principles through three architectural components: Multi-Vector Embeddings for context-dependent representation, Non-Collapsing Attention for parallel interpretation retention, and Contextual Identity Tracking (CIT) for maintaining $A\neq A$ across inference. We demonstrate NRR’s advantages through case studies in paradox handling, creative generation, and context-dependent reasoning. Crucially, we provide a minimal empirical validation on a synthetic context-shift task where an NRR-lite model achieves 90.9% out-of-distribution accuracy compared to 9.1% for standard architectures, demonstrating that ambiguity preservation enables structural generalization. NRR challenges the assumption that meaning must collapse to be useful, offering a foundation for AI systems capable of sophisticated ambiguity handling and creative reasoning. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.
© 2025 Kei Saito. Licensed under CC BY 4.0.
https://creativecommons.org/licenses/by/4.0/
Keywords: Non-resolution reasoning, contextual identity, ambiguity preservation, multi-vector embeddings, non-collapsing attention, structural generalization
1 Introduction
1.1 The Problem of Premature Semantic Collapse
Large language models (LLMs) have achieved remarkable success in natural language processing, demonstrating fluency in text generation, translation, and question-answering. However, these systems share a fundamental limitation: they over-commit to early resolution of ambiguity. This “premature semantic collapse bias” manifests as:
  • Binary decisions forced too quickly
  • Collapse of multiple valid interpretations into single outputs
  • Overconfident inference in uncertain contexts
  • Unnecessary computational branching and backtracking
  • Brittleness when handling paradoxes and self-referential structures
The root cause lies in classical identity assumptions embedded in neural architectures. Standard embedding layers assign each token a fixed vector (or context-dependent but ultimately collapsed representation). Attention mechanisms use softmax normalization, forcing probabilistic collapse to sum-to-one distributions. Generation objectives optimize for single “correct” outputs rather than maintaining interpretation flexibility.
Yet human cognition operates differently. Humans routinely maintain multiple interpretations simultaneously, suspend judgment when context is insufficient, and navigate contradictions without forcing resolution. This discrepancy suggests that AI’s struggles with ambiguity, paradox, creativity, and nuanced reasoning stem not from insufficient data or model scale, but from architectural assumptions about how meaning should be processed.
1.2 Contributions of This Work
This paper introduces Non-Resolution Reasoning (NRR), a computational framework that treats non-resolution—the deliberate maintenance of undecided states—as a valid and computationally useful reasoning mode. Our contributions are:
  1. 1.
    Conceptual Framework: We formalize three core principles (Non-Identity, Approximate Identity, Non-Resolution) that challenge classical identity assumptions in AI reasoning.
  2. 2.
    Architectural Components: We propose three implementable mechanisms—Multi-Vector Embeddings, Non-Collapsing Attention, and Contextual Identity Tracking (CIT)—adaptable to existing neural architectures.
  3. 3.
    Empirical Demonstration: We provide a minimal empirical validation on a synthetic context-shift task where NRR-lite achieves 90.9% out-of-distribution accuracy compared to 9.1% for standard architectures.
  4. 4.
    Theoretical Formalization: We provide mathematical foundations, including operators for non-identity ($\neq$) and non-resolution ($\aleph$), with explicit formalization of the $A\neq A\approx A$ principle.
  5. 5.
    Implementation Roadmap: We provide concrete benchmarking strategies and computational cost analysis to enable validation and extension.
NRR is not merely a refinement of existing methods but a paradigm shift: meaning does not need to collapse to be computationally useful.
2 The NRR Model: Foundational Principles
2.1 The $A\neq A\approx A$ Principle
At NRR’s core lies a rejection of classical identity assumptions. We summarize the intuition informally with the slogan $A\neq A\approx A$, which distinguishes contextual identity from structural similarity. Classical Logic: $A=A$ (Law of Identity). The same symbol always refers to the same entity. NRR Formalization:
$I(A,c_{1})\neq I(A,c_{2})\quad\text{and}\quad I(A,c_{1})\approx I(A,c_{2})$ (1)
where $I(A,c)$ denotes the contextual identity of symbol $A$ in context $c$, and $\approx$ is a graded similarity relation. Classical systems must either collapse to one meaning (lose information) or maintain ambiguity through probabilistic weighting (partial collapse). NRR maintains both as distinct identities while preserving their structural relationship.
2.2 Three Core Mechanisms
NRR operates through three principles:
(1) Non-Identity ($A\neq A$): The same symbol does not refer to the same entity across contexts. Identity is context-dependent, not universal.
(2) Approximate Identity ($A\approx A$): Entities share gradient-based similarity, not binary sameness. Similarity is continuous, allowing partial overlap without full identity.
(3) Non-Resolution: Conflicting interpretations can coexist without forced convergence. Ambiguity is a valid computational state, not an error.
3 Structural Limitations of Current AI
3.1 Why Existing AI Cannot Handle Ambiguity
Current models rely on reasoning frameworks derived from classical logic ($A=A$), probability maximization, and deterministic token prediction. These frameworks assume that meaning is stable across contexts and that a final resolution must be chosen. This leads to the collapse of polysemy (e.g., “light” must be illumination OR weight), suppression of contextual drift, and fragility in paradox handling.
3.2 Energy Efficiency Arguments
Retaining multiple interpretations appears computationally expensive. However, modern inference systems waste compute on premature resolution followed by backtracking and recomputing collapsed states when context shifts. We hypothesize that NRR can become energy-positive in settings such as long multi-turn dialogues with frequent context shifts and creative generation, where the upfront cost of maintaining parallel interpretations is offset by reduced backtracking.
4 NRR in Practice: Case Studies
We illustrate how NRR represents internal states in scenarios challenging for conventional architectures.
4.1 Paradox Handling: Internal Representation Matters
Problem: “This sentence is false.” Standard systems attempt binary truth evaluation, leading to contradiction or infinite loops. NRR maintains multiple contextual interpretations (truth_context, falsity_context, paradox_context) in parallel as a stable internal structure, allowing reasoning to continue without forced resolution. The paradox becomes data rather than error.
4.2 Creative Generation: Polysemy in Writing
Task: “Write a story about ‘light’.” Conventional AI disambiguates early (e.g., to “illumination”), producing single-layer output. NRR retains multiple meanings (illumination, weight, mood) simultaneously during generation, enabling multi-layered, poetic output where meanings are interwoven.
4.3 Context-Dependent Identity
Problem: Context shift in dialogue (e.g., “The bank is solid” $\to$ “ducks are swimming”). Conventional AI must backtrack and reprocess when the initial assumption (financial) proves wrong. NRR continuously reweights coexisting identities (financial vs. river) without backtracking, as both interpretations were maintained from the start.
5 Implementation Framework
This section presents architectural components for implementing NRR.
5.1 Multi-Vector Embeddings
Unlike standard embeddings that collapse to single vectors, NRR embeddings retain multiple contextual variants simultaneously.
$\mathcal{H}_{t}=\{\mathbf{h}_{t,1},\mathbf{h}_{t,2},\ldots,\mathbf{h}_{t,k}\}$ (2)
These can be trained via contrastive initialization or adversarial diversity loss to ensure variants capture distinct semantic dimensions.
5.2 Non-Collapsing Attention
Standard attention uses softmax normalization, forcing probabilistic collapse. NRR uses non-collapsing functions (e.g., sigmoid) to allow multiple positions to receive strong attention simultaneously without competition.
Algorithm: Non-Collapsing Attention Process
Input: Interpretation sets {H_t} for t=1 to n
Output: Updated interpretation sets {H’_t}
For each token t:
  For each interpretation i in {1...k}:
    1. Compute queries q, keys k, values v
    2. Compute scores: s_ij = q * k^T / sqrt(d)
    3. Apply Activation (No Softmax):
       alpha_ij = sigmoid(s_ij)
    4. Aggregate:
       h’_t,i = sum_j (alpha_ij * v_t,j)
  End
End
Figure 1: Conceptual logic for Non-Collapsing Attention. Unlike standard attention, scores are activated independently (e.g., via sigmoid) rather than competing via softmax.
Stability is maintained via gradient clipping and attention strength regularization.
5.3 Contextual Identity Tracking (CIT)
CIT explicitly tracks which interpretation belongs to which semantic context.
5.3.1 Context Detection
When semantic discontinuity exceeds a threshold:
$d(\mathbf{h}_{t},\mathbf{h}_{t-1})>\tau_{\text{context}}$ (3)
a new context ID is assigned. $\tau_{\text{context}}$ may be learned or fixed; in our minimal implementation, it is defined as the cosine-distance threshold separating intra-turn variability from inter-turn shifts.
5.4 Strategic Resolution: When and How to Collapse
While NRR treats non-resolution as computationally valid, practical systems must eventually act. We propose a resolution operator $\rho$ controlled by task affordances (classification vs. generation), contextual stability (context drift $\Delta c_{t}$), and interpretation dominance. Crucially, these policies are external to the NRR core, allowing the same representation to serve diverse tasks.
6 Empirical Validation
To demonstrate the $A\neq A$ principle, we implemented a minimal experiment validating the claim: maintaining contextual non-identity until sufficient context arrives improves generalization.
6.1 Task Design: Context-Switch Word Sense (CSWS)
We created a synthetic two-turn task where the identity of an ambiguous token (“bank”) is determined only after context arrives.
  • Turn 1 (Ambiguous): “The bank is {adj}.” (e.g., solid, stable)
  • Turn 2 (Context): Disambiguating cue (e.g., “investor” or “ducks”).
Task: Classify the sense of “bank” (FINANCIAL vs RIVER). Turn 1 provides insufficient information, forcing models to either commit early (risk error) or preserve ambiguity.
6.2 Model Comparison
  • Baseline Transformer: Standard single embedding per token.
  • NRR-lite: Maintains $k=2$ separate embeddings for the ambiguous token. Uses context gating in Turn 2 to resolve the embedding.
Note: NRR-lite is not a full implementation of NRR; it isolates the ambiguity-preservation component to demonstrate the causal effect of delayed resolution on generalization.
6.3 Results
Table 1: Minimal Experiment Results: Context-Switch Word Sense (CSWS) Task
Model ID Acc OOD Acc $H(g)$ Turn 1 $H(g)$ Turn 2
Baseline 1.000 0.091
NRR-lite 0.920 0.909 0.693 0.692
Figure 2: Minimal Experiment Results. Left: Ambiguity preservation measured by gate entropy $H(g)$. NRR-lite maintains high entropy at Turn 1 (no context). Right: Generalization performance. NRR-lite achieves 90.9% OOD accuracy while Baseline fails at 9.1%, demonstrating that ambiguity preservation enables structural rather than surface-level learning.
Key Findings:
  1. 1.
    Baseline Failure (9.1% OOD): The baseline memorized specific cue words but failed to generalize to unseen cues.
  2. 2.
    NRR Success (90.9% OOD): By maintaining ambiguity at Turn 1 and resolving based on Turn 2 context structure, NRR-lite learned a transferable representation.
  3. 3.
    Ambiguity Preservation: NRR-lite maintained maximum entropy ($\approx\log 2$) at Turn 1, confirming it did not commit prematurely.
This experiment validates that maintaining contextual non-identity ($A\neq A$) enables structural generalization where standard architectures fail.
7 Future Evaluation Framework
Comprehensive evaluation requires scaling beyond the minimal experiment. We propose benchmarks for:
  • Ambiguity Retention: Measuring entropy of interpretations under polysemy.
  • Context-Shift Adaptation: Measuring cost of adaptation vs. backtracking.
  • Creative Generation: Human evaluation of semantic richness.
  • Paradox Stability: Testing system stability under self-reference.
8 Conclusion
For readers interested in the theoretical underpinnings and philosophical framework of the proposed Non-Resolution Reasoning mechanism, please refer to our companion paper [34]. This paper introduced Non-Resolution Reasoning (NRR), a framework challenging classical identity assumptions in AI. We demonstrated through theory and minimal experimentation that preserving ambiguity is not a cost but a capability. Our experiment showed that delaying resolution allows for 90.9% generalization accuracy in ambiguous contexts where standard models fail (9.1%). NRR challenges the assumption that meaning must collapse to be useful. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.
Acknowledgments
The author gratefully acknowledges the support of large language models—including ChatGPT (OpenAI), Claude (Anthropic), and Gemini (Google)—for their assistance in linguistic refinement, LaTeX formatting, and general proofreading support during the preparation of this manuscript. All concepts, arguments, and conclusions remain solely the responsibility of the author.