### v8

# NRR-Core: Non-Resolution Reasoning as a Computational Framework for Contextual Identity and Ambiguity Preservation

Abstract
Current artificial intelligence systems, despite remarkable capabilities in text generation and pattern recognition, exhibit a fundamental architectural limitation: they resolve ambiguity prematurely. This premature semantic collapse—the tendency to collapse multiple valid interpretations into a single output—stems from classical identity assumptions embedded in standard neural architectures. We propose Non-Resolution Reasoning (NRR), a computational framework that treats ambiguity retention as a valid reasoning mode rather than a defect to be eliminated. NRR introduces three core principles: (1) Non-Identity ($A\neq A$)—the same symbol refers to different entities across contexts; (2) Approximate Identity ($A\approx A$)—entities share partial structural overlap without being identical; and (3) Non-Resolution—conflicting interpretations can coexist without forced convergence. We formalize these principles through three architectural components: Multi-Vector Embeddings for context-dependent representation, Non-Collapsing Attention for parallel interpretation retention, and Contextual Identity Tracking (CIT) for maintaining $A\neq A$ across inference. We provide functional verification of the Multi-Vector Embeddings component through a controlled synthetic task. We illustrate NRR’s potential through case studies in paradox handling, creative generation, and context-dependent reasoning. Crucially, we provide functional verification in a synthetic two-turn disambiguation task through a Turn 1 Entropy measurement demonstrating ambiguity preservation directly: NRR-lite maintains high entropy ($H=0.91$ bits, near-maximum $1.0$) at the ambiguous turn, while standard architectures collapse early ($H=0.15$ bits), demonstrating that NRR preserves interpretive flexibility until context arrives. NRR challenges the assumption that meaning must collapse to be useful, offering a foundation for AI systems capable of sophisticated ambiguity handling and creative reasoning. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.
Implementation: A reference implementation accompanying this work is available at https://github.com/kei-saito-research/nrr-core.
© 2025 Kei Saito. Licensed under CC BY 4.0.
https://creativecommons.org/licenses/by/4.0/
Keywords: Non-resolution reasoning, contextual identity, ambiguity preservation, multi-vector embeddings, non-collapsing attention, entropy measurement
1 Introduction
This paper is one module in an ongoing Non-Resolution Reasoning (NRR) research program, not a closed endpoint. Companion papers cover text-to-state mapping, implementation strategy, and cross-domain validation with distinct scopes. A series-level Program Map and update status will be maintained in repository documentation.
1.1 The Problem of Premature Semantic Collapse
Large language models (LLMs) have achieved remarkable success in natural language processing [1, 3], demonstrating fluency in text generation, translation, and question-answering. However, these systems share a fundamental limitation: they over-commit to early resolution of ambiguity. This “premature semantic collapse bias” manifests as:
  • Binary decisions forced too quickly
  • Collapse of multiple valid interpretations into single outputs
  • Overconfident inference in uncertain contexts
  • Unnecessary computational branching and backtracking
  • Brittleness when handling paradoxes and self-referential structures
The root cause lies in classical identity assumptions embedded in neural architectures. Standard embedding layers assign each token a fixed vector (or context-dependent but ultimately collapsed representation [2]). Attention mechanisms use softmax normalization [13], forcing probabilistic collapse to sum-to-one distributions. Generation objectives optimize for single “correct” outputs rather than maintaining interpretation flexibility. While recent decoding strategies introduce diversity at the output level [9, 10], the underlying representations still collapse internally.
Yet human cognition operates differently. Humans routinely maintain multiple interpretations simultaneously, suspend judgment when context is insufficient, and navigate contradictions without forcing resolution. This discrepancy suggests that AI’s struggles with ambiguity, paradox, creativity, and nuanced reasoning stem not from insufficient data or model scale, but from architectural assumptions about how meaning should be processed.
Several lines of prior work address related aspects of these limitations without targeting the core collapse problem. Efficient attention variants [8, 16] reduce computational cost but retain softmax normalization and its inherent collapse pressure. Mixture-of-experts architectures [7] maintain parallel computational pathways but route inputs to single experts, enforcing collapse at the routing level. Probabilistic programming languages [11] maintain distributions over hypotheses but typically aim for eventual convergence to point estimates. Neurosymbolic approaches [12] integrate logical and neural reasoning but inherit classical identity assumptions from their symbolic components. NRR differs from all of these by treating non-resolution—the sustained coexistence of conflicting interpretations—as a first-class computational primitive.
1.2 Contributions of This Work
This paper introduces Non-Resolution Reasoning (NRR), a computational framework that treats non-resolution—the deliberate maintenance of undecided states—as a valid and computationally useful reasoning mode. Our contributions are:
  1. 1.
    Conceptual Framework: We formalize three core principles (Non-Identity, Approximate Identity, Non-Resolution) that challenge classical identity assumptions in AI reasoning.
  2. 2.
    Architectural Components: We propose three implementable mechanisms—Multi-Vector Embeddings, Non-Collapsing Attention, and Contextual Identity Tracking (CIT)—adaptable to existing neural architectures.
  3. 3.
    Functional Verification: We validate ambiguity preservation directly via Turn 1 Entropy measurement, showing NRR-lite maintains $H=0.91$ bits (near-maximum) while the baseline collapses to $H=0.15$ bits, consistently across five random seeds.
  4. 4.
    Theoretical Formalization: We provide a mathematical outline, including a formal state space, a compact core operator set for non-collapsing transitions, and extensible metadata functions for context tracking, with explicit formalization of the $A\neq A\approx A$ principle.
  5. 5.
    Implementation Roadmap: We provide concrete benchmarking strategies and computational cost considerations to enable validation and extension.
NRR is not merely a refinement of existing methods but a paradigm shift: meaning does not need to collapse to be computationally useful.
2 The NRR Model: Foundational Principles
2.1 The $A\neq A\approx A$ Principle
At NRR’s core lies a rejection of classical identity assumptions. We summarize the intuition informally with the slogan $A\neq A\approx A$, which distinguishes contextual identity from structural similarity. Classical Logic: $A=A$ (Law of Identity). The same symbol always refers to the same entity. NRR Formalization:
$I(A,c_{1})\neq I(A,c_{2})\quad\text{and}\quad I(A,c_{1})\approx I(A,c_{2})$ (1)
where $I(A,c)$ denotes the contextual identity of symbol $A$ in context $c$, and $\approx$ is a graded similarity relation. Classical systems must either collapse to one meaning (lose information) or maintain ambiguity through probabilistic weighting (partial collapse). NRR maintains both as distinct identities while preserving their structural relationship.
2.2 Three Core Mechanisms
NRR operates through three principles:
(1) Non-Identity ($A\neq A$): The same symbol does not refer to the same entity across contexts. Identity is context-dependent, not universal.
(2) Approximate Identity ($A\approx A$): Entities share gradient-based similarity, not binary sameness. Similarity is continuous, allowing partial overlap without full identity.
(3) Non-Resolution: Conflicting interpretations can coexist without forced convergence. Ambiguity is a valid computational state, not an error.
3 Structural Limitations of Current AI
3.1 Why Existing AI Cannot Handle Ambiguity
Current models rely on reasoning frameworks derived from classical logic ($A=A$), probability maximization, and deterministic token prediction. These frameworks assume that meaning is stable across contexts and that a final resolution must be chosen. This leads to the collapse of polysemy (e.g., “light” must be illumination OR weight), suppression of contextual drift, and fragility in paradox handling.
3.2 Energy Efficiency Arguments
Retaining multiple interpretations appears computationally expensive. However, modern inference systems waste compute on premature resolution followed by backtracking and recomputing collapsed states when context shifts. We hypothesize that NRR can become energy-positive in settings such as long multi-turn dialogues with frequent context shifts and creative generation, where the upfront cost of maintaining parallel interpretations is offset by reduced backtracking.
4 NRR in Practice: Case Studies
We illustrate how NRR would represent internal states in scenarios challenging for conventional architectures. These are conceptual illustrations; empirical validation is provided in Section 6.
4.1 Paradox Handling: Internal Representation Matters
Problem: “This sentence is false.” Standard systems attempt binary truth evaluation, leading to contradiction or infinite loops. NRR would maintain multiple contextual interpretations (truth_context, falsity_context, paradox_context) in parallel as a stable internal structure, allowing reasoning to continue without forced resolution. The paradox becomes data rather than error.
4.2 Creative Generation: Polysemy in Writing
Task: “Write a story about ‘light’.” Conventional AI disambiguates early (e.g., to “illumination”), producing single-layer output. NRR would retain multiple meanings (illumination, weight, mood) simultaneously during generation, enabling multi-layered, poetic output where meanings are interwoven.
4.3 Context-Dependent Identity
Problem: Context shift in dialogue (e.g., “The bank is solid” $\to$ “ducks are swimming”). Conventional AI must backtrack and reprocess when the initial assumption (financial) proves wrong. NRR would continuously reweight coexisting identities (financial vs. river) without backtracking, as both interpretations were maintained from the start.
5 Implementation Framework
This section presents architectural components for implementing NRR. While related to prior work on multi-prototype embeddings [4, 5, 6] and non-softmax attention [14, 15], the mechanisms in NRR are unified toward explicit ambiguity preservation rather than sense selection or sparsity.
5.1 Multi-Vector Embeddings
Unlike standard embeddings that collapse to single vectors, NRR embeddings are designed to retain multiple contextual variants simultaneously.
$\mathcal{H}_{t}=\{\mathbf{h}_{t,1},\mathbf{h}_{t,2},\ldots,\mathbf{h}_{t,k}\}$ (2)
These could be trained via contrastive initialization or adversarial diversity loss to ensure variants capture distinct semantic dimensions; we leave concrete training schemes to future empirical work.
5.2 Non-Collapsing Attention
Standard attention uses softmax normalization, forcing probabilistic collapse. NRR proposes non-collapsing activation functions (e.g., sigmoid or other non-competitive functions) to allow multiple positions to receive strong attention simultaneously without competition.
Algorithm: Non-Collapsing Attention Process
Input: Interpretation sets {H_t} for t=1 to n
Output: Updated interpretation sets {H’_t}
For each token t:
  For each interpretation i in {1...k}:
    1. Compute queries q, keys k, values v
    2. Compute scores: s_ij = q * k^T / sqrt(d)
    3. Apply Activation (No Softmax):
       alpha_ij = sigmoid(s_ij)
    4. Aggregate:
       h’_t,i = sum_j (alpha_ij * v_t,j)
  End
End
Figure 1: Conceptual logic for Non-Collapsing Attention. Unlike standard attention, scores are activated independently (e.g., via sigmoid) rather than competing via softmax.
Stability could be maintained through techniques such as layer normalization [18] and residual scaling [17], along with gradient clipping and attention strength regularization.
5.3 Contextual Identity Tracking (CIT)
CIT explicitly tracks which interpretation belongs to which semantic context.
5.3.1 Context Detection
When semantic discontinuity exceeds a threshold:
$d(\mathbf{h}_{t},\mathbf{h}_{t-1})>\tau_{\text{context}}$ (3)
a new context ID is assigned. $\tau_{\text{context}}$ may be learned or fixed; in a minimal implementation, it could be defined as the cosine-distance threshold separating intra-turn variability from inter-turn shifts.
5.4 Strategic Resolution: When and How to Collapse
While NRR treats non-resolution as computationally valid, practical systems must eventually act. We propose a resolution operator controlled by task affordances (classification vs. generation), contextual stability (context drift $\Delta c_{t}$), and interpretation dominance. Crucially, these policies are external to the NRR core, allowing the same representation to serve diverse tasks.
5.5 Protocol Extensions: State Space and Operators
The architectural components above (MVE, NCA, CIT) can be unified under a formal state space that enables systematic extension. This section provides the minimal specification for researchers implementing NRR-based systems.
5.5.1 State Space Formalization
The non-collapsing state is represented as:
$\mathcal{S}=\bigl((v_{i},c_{i},w_{i},m_{i})\bigr)_{i=1}^{n}$ (4)
where $v_{i}\in\mathbb{R}^{d}$ is a semantic vector, $c_{i}$ is a context identifier, $w_{i}\in[0,1]$ is an activation weight, and $m_{i}$ is optional metadata. Crucially, the weights $w_{i}$ are not normalized to sum to unity—multiple interpretations can maintain high activation simultaneously without competing for probability mass. This is the formal expression of non-collapse.
For measurement purposes, we associate a discrete distribution over interpretations:
$p(i\mid\mathcal{S})=\frac{w_{i}}{\sum_{j}w_{j}}$ (5)
and define $H(\mathcal{S})$ as the Shannon entropy of this distribution:
$H(\mathcal{S})=-\sum_{i=1}^{n}p(i\mid\mathcal{S})\log_{2}p(i\mid\mathcal{S})$ (6)
This normalization is used only for evaluation; the internal NRR dynamics do not require $w_{i}$ to sum to unity.
5.5.2 Interference Operators
We introduce a candidate family of interference operators that act on $\mathcal{S}$. The set is extensible rather than fixed; practical implementations may use a minimal subset. We define an NRR interference operator as a transformation $T$ on $\mathcal{S}$ whose intended property is approximate information preservation, i.e., our design goal is that
$H(T(\mathcal{S}))\geq H(\mathcal{S})-\epsilon$ (7)
for small $\epsilon$. Table 1 specifies the operators and their roles; verifying and refining their information-preservation guarantees is left to future empirical and theoretical work. Notation note: Later implementation papers distinguish state-level and item-level operators. In particular, IME/Universal use item-level updates (often denoted $\sigma_{\text{item}},\delta_{\text{item}}$), while this table lists the broader state-level candidate family. One operator, CPP Integration ($\kappa$), implements the Contradiction-Preservation Principle: contradictory hypotheses are retained rather than eliminated.
Operator Symbol Function
State-level Calibration $\sigma$ Uniform scaling/check operator (may be entropy-invariant)
Label-Free Abstraction $\alpha$ Projects to geometric relations without labels
Positioning $\rho$ Assigns contextual coordinates (implements CIT)
Invariance $\iota$ Identifies structure stable across contexts
Dampening $\delta$ Reduces activation differentials
Deferred Resolution $\tau$ Postpones projection to output boundary
CPP Integration $\kappa$ Preserves contradictory states with conflict tags
Persistence $\pi$ Maintains state across turns with decay
Table 1: NRR interference operators. Each is intended to transform $\mathcal{S}$ in ways that avoid collapse.
One possible pipeline: $\mathcal{S}^{\prime}=\pi(\kappa(\delta(\rho(\sigma(\mathcal{S}))),\mathcal{S}_{\text{new}}))$—stripping biases, positioning, dampening, integrating new information, and persisting.
Note on operator constraints: While NRR permits non-resolution as a valid state, not all transformations preserve this property. Operators must satisfy design principles (e.g., relative structure preservation) to avoid collapsing the very ambiguity NRR maintains. The formal specification of these constraints is addressed in subsequent work.
5.5.3 Non-Destructive Projection
To produce singular output while preserving internal state, NRR uses non-destructive projection:
$\Pi(\mathcal{S})=g(v_{i^{*}}),\quad i^{*}=\arg\max_{i}w_{i},\quad\mathcal{S}^{\prime}=\mathcal{S}$ (8)
where $g:\mathbb{R}^{d}\to\mathcal{Y}$ maps the semantic vector to the output space (e.g., token vocabulary). The system outputs $g(v_{i^{*}})$ but retains the full state $\mathcal{S}$. This contrasts with collapse-based systems where output destroys alternatives. The user receives a singular answer; the system preserves interpretive depth for subsequent reasoning.
6 Functional Verification
Scope of this section: This section provides a controlled synthetic evaluation demonstrating that the proposed architecture behaves as intended: it preserves ambiguity before contextual cues arrive and resolves it afterward. This is not a comparison against modern LLMs but a minimal functional verification (proof-of-concept). Comparative evaluation against production systems is reserved for future work.
To demonstrate the $A\neq A$ principle, we designed a verification procedure that directly measures ambiguity preservation: does the model maintain interpretive uncertainty when context is insufficient?
6.1 Task Design: Turn 1 Entropy Measurement
We created a synthetic two-turn disambiguation task where the identity of an ambiguous token (“bank”) is determined only after context arrives.
  • Turn 1 (Ambiguous): “The bank is {adj}.” (e.g., solid, stable, old, new)
  • Turn 2 (Context): Disambiguating cue (e.g., “investor” $\to$ FINANCIAL, “ducks” $\to$ RIVER).
Key Insight: Rather than measuring classification accuracy (which depends on label distribution), we directly measure entropy of the output distribution at Turn 1—before any disambiguating context arrives.
$H(p)=-\sum_{i}p_{i}\log_{2}p_{i}$ (9)
For a binary classification:
  • $H_{\max}=\log_{2}2=1.0$ bit (perfect ambiguity: $p=[0.5,0.5]$)
  • $H_{\min}=0$ (complete certainty: $p=[1,0]$ or $[0,1]$)
A model that preserves ambiguity should maintain high entropy at Turn 1; a model that collapses early will show low entropy.
6.2 Model Comparison
  • Baseline: Standard single embedding per token, 32-dimensional embeddings, 2-layer MLP classifier.
  • NRR-lite: Implements the MVE principle in minimal form, maintaining $k=2$ separate embeddings for the ambiguous token (“bank”). A context-dependent gate selects the appropriate embedding based on Turn 2 context. When Turn 2 is neutral (no context), the gate remains balanced.
Both models were trained on 1,000 samples with balanced labels (50% FINANCIAL, 50% RIVER) for 100 epochs. All reported entropy values and statistics are computed from the trained models over 5 independent random seeds.
6.3 Results
Table 2: Turn 1 Entropy Verification Results (5 random seeds)
Model Turn 1 Entropy $H$ (bits) Gate Entropy Context Accuracy
Baseline $0.15\pm 0.13$ 100%
NRR-lite $\mathbf{0.91\pm 0.04}$ $1.00$ 100%
Maximum $H$ $1.00$ $1.00$
Figure 2: Turn 1 Entropy Verification Results (aggregated across 5 seeds). Left: Entropy at Turn 1 (no context). NRR-lite maintains high entropy ($H\approx 0.91$ bits), near the theoretical maximum, while Baseline collapses ($H\approx 0.15$ bits). Center: Distribution of $P(\text{FINANCIAL})$ at Turn 1 across all test samples (500 total: 5 seeds × 100 samples each). NRR-lite centers around 0.5 (ambiguous); Baseline clusters near 1.0 (committed to financial interpretation). Right: Per-seed Turn 1 Entropy breakdown across all 5 random seeds, demonstrating consistent preservation by NRR-lite (blue, $H\approx 0.9$ bits across seeds) and variable collapse by Baseline (red, ranging from 0.01 to 0.37 bits depending on initialization).
Key Findings:
  1. 1.
    Baseline Collapse ($H=0.15$ bits): The baseline commits to one interpretation at Turn 1, even without disambiguating context. This premature collapse is the architectural bias NRR addresses.
  2. 2.
    NRR-lite Preservation ($H=0.91$ bits): NRR-lite maintains near-maximum entropy at Turn 1, confirming that multiple interpretations coexist until context arrives.
  3. 3.
    Gate Behavior: NRR-lite’s context gate achieves $H=1.0$ bit (perfect balance) when no context is provided, shifting appropriately when Turn 2 arrives.
  4. 4.
    Resolution Capability: Both models achieve 100% accuracy with context, demonstrating that ambiguity preservation does not impair eventual resolution.
  5. 5.
    Robustness: Results replicate across 5 random seeds (paired $t$-test: $t(4)=11.9$, $p<0.001$).
  6. 6.
    Baseline Instability: Baseline’s high standard deviation ($\text{SD}=0.13$, 87% of mean) reflects architectural instability across random initializations, with entropy ranging from 0.01 to 0.37 bits depending on seed. This contrasts sharply with NRR-lite’s consistent performance ($\text{SD}=0.04$, 4% of mean), demonstrating robust ambiguity preservation independent of initialization.
This verification provides evidence that NRR-lite implements the $A\neq A$ principle: the same symbol (“bank”) maintains distinct potential identities until context determines which applies.
7 Future Evaluation Framework
Comprehensive evaluation requires scaling beyond this minimal verification. We propose benchmarks for:
  • Ambiguity Retention: Measuring entropy of interpretations under polysemy.
  • Context-Shift Adaptation: Measuring cost of adaptation vs. backtracking.
  • Creative Generation: Human evaluation of semantic richness.
  • Paradox Stability: Testing system stability under self-reference.
8 Conclusion
This paper introduced Non-Resolution Reasoning (NRR), a framework challenging classical identity assumptions in AI. In our controlled setting, we showed that preserving ambiguity can function as a capability rather than a cost. Our Turn 1 Entropy verification showed that NRR-lite maintains high interpretive entropy ($H=0.91$ bits) when context is absent, while standard architectures collapse prematurely ($H=0.15$ bits)—yet both achieve perfect accuracy when context arrives. NRR challenges the assumption that meaning must collapse to be useful. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.
Acknowledgments
The author gratefully acknowledges the support of large language models—including Claude (Anthropic), ChatGPT and Codex (OpenAI), and Gemini (Google)—for their assistance in linguistic refinement, LaTeX formatting, and general proofreading support during the preparation of this manuscript. All concepts, arguments, and conclusions remain solely the responsibility of the author.